# Is Chain-of-Thought Reasoning of LLMs a Mirage? A Data Distribution Lens

### v3

@@ Abstract
Chain-of-Thought (CoT) prompting has been shown to improve Large Language Model (LLM) performance on various tasks. With this approach, LLMs appear to produce human-like reasoning steps before providing answers (a.k.a., CoT reasoning), which often leads to the perception that they engage in deliberate inferential processes. However, some initial findings suggest that CoT reasoning may be more superficial than it appears, motivating us to explore further. In this paper, we study CoT reasoning via a data distribution lens and investigate if CoT reasoning reflects a structured inductive bias learned from in-distribution data, allowing the model to conditionally generate reasoning paths that approximate those seen during training. Thus, its effectiveness is fundamentally bounded by the degree of distribution discrepancy between the training data and the test queries. With this lens, we dissect CoT reasoning via three dimensions: task, length, and format. To investigate each dimension, we design DataAlchemy, an isolated and controlled environment to train LLMs from scratch and systematically probe them under various distribution conditions. Our results reveal that CoT reasoning is a brittle mirage that vanishes when it is pushed beyond training distributions. This work offers a deeper understanding of why and when CoT reasoning fails, emphasizing the ongoing challenge of achieving genuine and generalizable reasoning. Our code is available at GitHub: https://github.com/ChengshuaiZhao0/DataAlchemy.
@@ 1 Introduction
Figure 1: The data perspective lens. CoT reasoning’s effectiveness is fundamentally bounded by the degree of distribution discrepancy between the training data and the test queries. Guided by this lens, we dissect CoT reasoning via three dimensions: task, length, and format.
@@
Recent years have witnessed Large Language Models’ (LLMs) dominant role in various domains (Zhao et al., 2023; Li et al., 2025b; Zhao et al., 2025; Ting et al., 2025) through versatile prompting techniques (Wei et al., 2022; Yao et al., 2023; Kojima et al., 2022). Among these, Chain-of-Thought (CoT) prompting (Wei et al., 2022) has emerged as a prominent method for eliciting structured reasoning from LLMs (a.k.a., CoT reasoning). By appending a simple cue such as “Let’s think step by step,” LLMs decompose complex problems into intermediate steps, producing outputs that resemble human-like reasoning. It has been shown to be effective in tasks requiring logical inference(Xu et al., 2024), mathematical problem solving (Imani et al., 2023), and commonsense reasoning (Wei et al., 2022). The empirical successes of CoT reasoning lead to the perception that LLMs engage in deliberate inferential processes (Yu et al., 2023; Zhang et al., 2024a; Ling et al., 2023; Zhang et al., 2024c).
However, a closer examination reveals inconsistencies that challenge this optimistic view. Consider this straightforward question: “The day the US was established is in a leap year or a normal year?” When prompted with the CoT prefix, the modern LLM Gemini responded: “The United States was established in 1776. 1776 is divisible by 4, but it’s not a century year, so it’s a leap year. Therefore, the day the US was established was in a normal year.” This response exemplifies a concerning pattern: the model correctly recites the leap year rule and articulates intermediate reasoning steps, yet produces a logically inconsistent conclusion (i.e., asserting 1776 is both a leap year and a normal year). Such inconsistencies suggest that there is a distinction between human-like inference and CoT reasoning.
An expanding body of analyses reveals that LLMs tend to rely on surface-level semantics and clues rather than logical procedures (Stechly et al., 2024; Kambhampati, 2024; Chen et al., 2025b; Lanham et al., 2023). LLMs construct superficial chains of logic based on learned token associations, often failing on tasks that deviate from commonsense heuristics or familiar templates (Tang et al., 2023). In the reasoning process, performance degrades sharply when irrelevant clauses are introduced, which indicates that models cannot grasp the underlying logic (Mirzadeh et al., 2024). This fragility becomes even more apparent when models are tested on more complex tasks, where they frequently produce incoherent solutions and fail to follow consistent reasoning paths (Shojaee et al., 2025). Collectively, these pioneering works deepen the skepticism surrounding the true nature of CoT reasoning.
In light of this line of research, we question the CoT reasoning by proposing an alternative lens through data distribution and further investigating why and when it fails. We hypothesize that CoT reasoning reflects a structured inductive bias learned from in-distribution data, allowing the model to conditionally generate reasoning paths that approximate those seen during training. As such, its effectiveness is inherently limited by the nature and extent of the distribution discrepancy between training data and the test queries. Guided by this data distribution lens, we dissect CoT reasoning via three dimensions: (i) task—To what extent CoT reasoning can handle tasks that involve transformations or previously unseen task structures. (2) length—how CoT reasoning generalizes to chains with length different from that of training data; and (3) format—how sensitive CoT reasoning is to surface-level query form variations. To evaluate each aspect, we introduce DataAlchemy, a controlled and isolated experiment that allows us to train LLMs from scratch and systematically probe them under various distribution shifts.
@@
Our findings reveal that CoT reasoning works effectively when applied to in-distribution or near in-distribution data but becomes fragile and prone to failure even under moderate distribution shifts. In some cases, LLMs generate fluent yet logically inconsistent reasoning steps. The results suggest that what appears to be structured reasoning can be a mirage, emerging from memorized or interpolated patterns in the training data rather than logical inference. These insights carry important implications for both practitioners and researchers. For practitioners, our results highlight the risk of relying on CoT as a plug-and-play solution for reasoning tasks and caution against equating CoT-style output with human thinking. For researchers, the results underscore the ongoing challenge of achieving reasoning that is both faithful and generalizable, motivating the need to develop models that can move beyond surface-level pattern recognition to exhibit deeper inferential competence. Our contributions are summarized as follows:
@@
    Novel perspective. We propose a data distribution lens for CoT reasoning, illuminating that its effectiveness stems from structured inductive biases learned from in-distribution training data. This framework provides a principled lens for understanding why and when CoT reasoning succeeds or fails.
@@
    Controlled environment. We introduce DataAlchemy, an isolated experimental framework that enables training LLMs from scratch and systematically probing CoT reasoning. This controlled setting allows us to isolate and analyze the effects of distribution shifts on CoT reasoning without interference from complex patterns learned during large-scale pre-training.
@@
    Empirical validation. We conduct systematic empirical validation across three critical dimensions—task, length, and format. Our experiments demonstrate that CoT reasoning exhibits sharp performance degradation under distribution shifts, revealing that seemingly coherent reasoning masks shallow pattern replication.
@@
    Real-world implication. This work reframes the understanding of contemporary LLMs’ reasoning capabilities and emphasizes the risk of over-reliance on COT reasoning as a universal problem-solving paradigm. It underscores the necessity for proper evaluation methods and the development of LLMs that possess authentic and generalizable reasoning capabilities.
@@ 2.1 LLM Prompting and CoT
Chain-of-Thought (CoT) prompting revolutionized how we elicit reasoning from Large Language Models by decomposing complex problems into intermediate steps (Wei et al., 2022). By augmenting few-shot exemplars with reasoning chains, CoT showed substantial performance gains on various tasks (Xu et al., 2024; Imani et al., 2023; Wei et al., 2022). Building on this, several variants emerged. Zero-shot CoT triggers reasoning without exemplars using instructional prompts (Kojima et al., 2022), and self-consistency enhances performance via majority voting over sampled chains (Wang et al., 2023). To reduce manual effort, Auto-CoT generates CoT exemplars using the models themselves (Zhang et al., 2023). Beyond linear chains, Tree-of-Thought (ToT) frames CoT as a tree search over partial reasoning paths (Yao et al., 2023), enabling lookahead and backtracking. SymbCoT combines symbolic reasoning with CoT by converting problems into formal representations (Xu et al., 2024). Recent work increasingly integrates CoT into the LLM inference process, generating long-form CoTs (Jaech et al., 2024; Team, 2024; Guo et al., 2025; Team et al., 2025). This enables flexible strategies like mistake correction, step decomposition, reflection, and alternative reasoning paths (Yeo et al., 2025; Chen et al., 2025a). The success of prompting techniques and long-form CoTs has led many to view them as evidence of emergent, human-like reasoning in LLMs. In this work, we challenge that viewpoint by adopting a data-centric perspective and demonstrating that CoT behavior arises largely from pattern matching over training distributions.
@@ 2.2 Discussion on Illusion of LLM Reasoning
While Chain-of-Thought prompting has led to impressive gains on complex reasoning tasks, a growing body of work has started questioning the nature of these gains (Stechly et al., 2024, 2025; Kambhampati et al., 2025). One major line of research highlights the fragility of CoT reasoning. Minor and semantically irrelevant perturbations such as distractor phrases or altered symbolic forms can cause significant performance drops in state-of-the-art models (Mirzadeh et al., 2024; Tang et al., 2023). Models often incorporate such irrelevant details into their reasoning, revealing a lack of sensitivity to salient information. Other studies show that models prioritize the surface form of reasoning over logical soundness; in some cases, longer but flawed reasoning paths yield better final answers than shorter, correct ones (Bentham et al., 2024). Similarly, performance does not scale with problem complexity as expected—models may overthink easy problems and give up on harder ones (Shojaee et al., 2025). Another critical concern is the faithfulness of the reasoning process. Intervention-based studies reveal that final answers often remain unchanged even when intermediate steps are falsified or omitted (Lanham et al., 2023), a phenomenon dubbed the illusion of transparency (Bentham et al., 2024; Chen et al., 2025b). Together, these findings suggest that LLMs are not principled reasoners but rather sophisticated simulators of reasoning-like text. However, a systematic understanding of why and when CoT reasoning fails is still a mystery.
@@ 2.3 OOD Generalization of LLMs
Out-of-distribution (OOD) generalization, where test inputs differ from training data, remains a key challenge in machine learning, particularly for large language models (LLMs)(Yang et al., 2024, 2023; Budnikov et al., 2025; Zhang et al., 2024b). Recent studies show that LLMs prompted to learn novel functions often revert to similar functions encountered during pretraining (Wang et al., 2024; Garg et al., 2022). Likewise, LLM generalization frequently depends on mapping new problems onto familiar compositional structures (Song et al., 2025). CoT prompting improves OOD generalization (Wei et al., 2022), with early work demonstrating length generalization for multi-step problems beyond training distributions (Yao et al., 2025; Shen et al., 2025). However, this ability is not inherent to CoT and heavily depends on model architecture and training setups. For instance, strong generalization in arithmetic tasks was achieved only when algorithmic structures were encoded into positional encodings (Cho et al., 2024). Similarly, finer-grained CoT demonstrations during training boost OOD performance, highlighting the importance of data granularity (Wang et al., 2025a). Theoretical and empirical evidence shows that CoT generalizes well only when test inputs share latent structures with training data; otherwise, performance declines sharply (Wang et al., 2025b; Li et al., 2025a). Despite its promise, CoT still struggles with genuinely novel tasks or formats. In the light of these brilliant findings, we propose rethinking CoT reasoning through a data distribution lens: decomposing CoT into task, length, and format generalization, and systematically investigating each in a controlled setting.
Figure 2: Framework of DataAlchemy. It creates an isolated and controlled environment to train LLMs from scratch and probe the task, length, and format generalization.
3 The Data Distribution Lens
We propose a fundamental reframing to understand what CoT actually represents. We hypothesize that the underlying mechanism is better understood through the lens of data distribution: rather than executing explicit reasoning procedures, CoT operates as a pattern-matching process that interpolates and extrapolates from the statistical regularities present in its training distribution. Specifically, we posit that CoT’s success stems not from a model’s inherent reasoning capacity, but from its ability to generalize conditionally to out-of-distribution (OOD) test cases that are structurally similar to in-distribution exemplars.
@@ 3 The Data Distribution Lens
To formalize this view, we model CoT prompting as a conditional generation process constrained by the distributional properties of the training data. Let $\mathcal{D}_{\text{train}}$ denote the training distribution over input-output pairs $(x,y)$, where $x$ represents a reasoning problem and $y$ denotes the solution sequence (including intermediate reasoning steps). The model learns an approximation $f_{\theta}(x)\approx y$ by minimizing empirical risk over samples drawn from $\mathcal{D}_{\text{train}}$.
Let the expected training risk be defined as:
$R_{\text{train}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{train}}}[\ell(f_{\theta}(x),y)],$ (1)
@@
where $\ell$ is a task-specific loss function (e.g., cross-entropy, token-level accuracy). At inference time, given a test input $a_{\text{test}}$ sampled from a potentially different distribution $\mathcal{D}_{\text{test}}$, the model generates a response $y_{\text{test}}$ conditioned on patterns learned from $\mathcal{D}_{\text{train}}$. The corresponding expected test risk is:
$R_{\text{test}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}[\ell(f_{\theta}(x),y)].$ (2)
The degree to which the model generalizes from $\mathcal{D}_{\text{train}}$ to $\mathcal{D}_{\text{test}}$ is governed by the distributional discrepancy between the two, which we quantify using divergence measures:
Definition 3.1 (Distributional Discrepancy).
@@
Given training distribution $\mathcal{D}_{\text{train}}$ and test distribution $\mathcal{D}_{\text{test}}$, the distributional discrepancy is defined as:
$\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})=\mathcal{H}(\mathcal{D}_{\text{train}}\parallel\mathcal{D}_{\text{test}})$ (3)
where $\mathcal{H}(\cdot\parallel\cdot)$ is a divergence measure (e.g., KL divergence, Wasserstein distance) that quantifies the statistical distance between the two distributions.
Theorem 3.1 (CoT Generalization Bound).
Let $f_{\theta}$ denote a model trained on $\mathcal{D}_{\text{train}}$ with expected training risk $R_{\text{train}}(f_{\theta})$. For a test distribution $\mathcal{D}_{\text{test}}$, the expected test risk $R_{\text{test}}(f_{\theta})$ is bounded by:
@@
$R_{\text{test}}(f_{\theta})\leq R_{\text{train}}(f_{\theta})+\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right)$ (4)
where $\Lambda>0$ is a Lipschitz constant that depends on the model architecture and task complexity, $n$ is the training sample size, and the bound holds with probability $1-\delta$, where $\delta$ is the failure probability.
The proof is provided in Appendix A.1
Building on this data distribution perspective, we identify three critical dimensions along which distributional shifts can occur, each revealing different aspects of CoT’s pattern-matching nature: ➊ Task generalization examines how well CoT transfers across different types of reasoning tasks. Novel tasks may have unique elements and underlying logical structure, which introduces distributional shifts that challenge the model’s ability to apply learned reasoning patterns. ➋ Length generalization investigates CoT’s robustness to reasoning chains of varying lengths. Since training data typically contains reasoning sequences within a certain length range, test cases requiring substantially longer or shorter reasoning chains represent a form of distributional shift along the sequence length dimension. This length discrepancy could result from the reasoning step or the text-dependent solution space. ➌ Format generalization explores how sensitive CoT is to variations in prompt formulation and structure. Due to various reasons (e.g., sophistical training data or diverse background of users), it is challenging for LLM practitioners to design a golden prompt to elicit knowledge suitable for the current case. Their detailed definition and implementation are given in subsequent sections.
Each dimension provides a unique lens for understanding the boundaries of CoT’s effectiveness and the mechanisms underlying its apparent reasoning capabilities. By systematically varying these dimensions in controlled experimental settings, we can empirically validate our hypothesis that CoT performance degrades predictably as distributional discrepancy increases, thereby revealing its fundamental nature as a pattern-matching rather than reasoning system.
4 DataAlchemy: An Isolated and Controlled Environment
To systematically investigate the influence of distributional shifts on CoT reasoning capabilities, we introduce DataAlchemy, a synthetic dataset framework designed for controlled experimentation. This environment enables us to train language models from scratch under precisely defined conditions, allowing for rigorous analysis of CoT behavior across different OOD scenarios. The overview is shown in Figure 2.

### v4

@@ Abstract
Chain-of-Thought (CoT) prompting has been shown to be effective in eliciting structured reasoning (i.e., CoT reasoning) from large language models (LLMs). Regardless of its popularity, recent studies expose its failures in some reasoning tasks, raising fundamental questions about the nature of CoT reasoning. In this work, we propose a data distribution lens to understand when and why CoT reasoning succeeds or fails. We hypothesize that CoT reasoning reflects a structured inductive bias learned from in-distribution data, enabling models to conditionally generate reasoning trajectories that approximate those observed during training. As such, the effectiveness of CoT reasoning is fundamentally governed by the nature and degree of distribution discrepancy between training data and test queries. Guided by this lens, we dissect CoT reasoning via three dimensions: task, length, and format. To test the hypothesis, we introduce DataAlchemy, an abstract and fully controllable environment that trains LLMs from scratch and systematically probes them under various distribution conditions. Through rigorous controlled experiments, we reveal that CoT reasoning is a brittle mirage when it is pushed beyond training distributions, emphasizing the ongoing challenge of achieving genuine and generalizable reasoning. Our code is available at GitHub: https://github.com/ChengshuaiZhao0/DataAlchemy.
@@ 1 Introduction
Figure 1: The data perspective lens. CoT reasoning’s effectiveness is fundamentally bounded by the degree of distribution discrepancy introduced by task, length, and format between the training data and the test queries.
@@
Chain-of-Thought (CoT) prompting wei2022chain has emerged as a prominent method for eliciting structured reasoning from LLMs (a.k.a., CoT reasoning). By appending a simple cue such as “Let’s think step by step”, LLMs decompose complex problems into intermediate steps, producing outputs that resemble human-like reasoning. It has been shown to be effective in tasks requiring logical inference xu2024faithful, mathematical problem solving imani2023mathprompter, and commonsense reasoning wei2022chain. The empirical success led to CoT reasoning being seen as a promising direction towards artificial general intelligence.
However, some pioneering studies have revealed failures that challenge this optimistic view mirzadehgsm. stechly2024chain demonstrate that LLMs fail to generalize in planning tasks, revealing a deficiency in true algorithmic reasoning. shojaee2025illusion find that reasoning models experience an accuracy collapse in puzzle-solving once task complexity exceeds a critical threshold. sun2025omega demonstrate that LLMs struggle to solve complex mathematical reasoning problem, failing to integrate or adapt learned skills to novel or creative tasks.
Considering the above opposing opinions, there clearly lacks an indisputable lens to understand why and when CoT reasoning succeeds or fails. Current evaluation approaches have intrinsic limitations that prevent them from answering the Why and When questions: (i) Narrowly defined settings. Existing frameworks focus on specific tasks and evaluate them using specific LLMs, thereby overlooking common structural patterns and characteristics. (ii) Data entanglement. Most evaluations are conducted in real-world scenarios, where the complexity precludes fully controlled experiments to isolate fine-grained factors. (iii) Data leakage. Pre-trained LLMs suffer from data leakage and benchmark containment problems, undermining the effectiveness and validity of evaluations.
In this work, we study CoT reasoning by introducing a data distribution lens. Specifically, we hypothesize that CoT reasoning reflects a structured inductive bias learned from in-distribution data, enabling models to conditionally generate reasoning trajectories that approximate those observed during training. As such, the effectiveness of CoT reasoning is fundamentally governed by the nature and degree of distribution discrepancy between training data and test queries. Guided by this lens, we revisit existing NLP tasks and identify three primary axes along which distribution shifts may occur: task (i.e., unseen task structures), length (i.e., different text lengths and reasoning lengths), and format (i.e., query format variant).
To tackle the issue of evaluations and validate our hypothesis, we further introduce DataAlchemy, an abstract, controllable, and clean environment. DataAlchemy provides an abstract representation system that distills various real-world NLP tasks into key components: atoms (i.e., token space), elements (i.e., text space), and transformations (i.e., operation space). By varying these components, we curate data that exhibits various distribution discrepancies, naturally achieving full and fine-grained control over the entire evaluation pipeline. Later, we train models from scratch to avoid data leakage and employ controlled experiments to rigorously test our hypotheses.
@@
Our findings reveal that CoT reasoning works effectively when applied to (near) in-distribution data, but becomes fragile and prone to failure even under moderate distribution shifts. In some cases, LLMs generate fluent yet logically inconsistent reasoning steps. The results suggest that what appears to be structured reasoning can be a mirage, emerging from memorized or interpolated patterns in the training data rather than logical inference. Our contributions can be summarized as follows:
@@
    Novel perspective. We propose a data distribution lens for CoT reasoning, revealing that its effectiveness arises from structured inductive biases learned from in-distribution data. This lens offers a principled foundation for understanding why and when CoT reasoning succeeds or fails.
@@
    Controllable environment. We develop an abstract, fully controllable, and clean environment—DataAlchemy that abstracts NLP tasks, enabling systematic analysis of CoT reasoning under distribution discrepancies. DataAlchemy can serve as a research platform for probing the intrinsic behavior of LLMs, and facilitating the discovery of scientific principles.
@@
    Rigorous investigation. Guided by data distribution lens, we dissect the CoT reasoning via three dimensions: task, length, and format. Later, we curate data that reflects fine-grained factors in each dimension and conduct controlled experiments to isolate and examine each factor.
@@
    General validity. We train and fine-tune hundreds of LLMs with varying sizes (from 62K to 14B), architectures (e.g., GPT, LLaMA, and Qwen), and temperatures (from 1e-5 to 10). The results consistently show that the effectiveness of CoT reasoning varies with the degree of distribution discrepancy, substantiating the generality of the proposed data distribution lens.
@@ 2.1 LLM Prompting and CoT
Chain-of-Thought (CoT) prompting improves large language model performance by eliciting intermediate reasoning steps for complex problems (wei2022chain). Extensions include zero-shot CoT (kojima2022large), self-consistency via sampling and voting (wang2023selfconsistency), and Auto-CoT, which automatically generates reasoning exemplars (zhang2023automatic). Beyond linear reasoning, Tree-of-Thought enables search over multiple reasoning paths (yao2023tree), while SymbCoT integrates symbolic representations into CoT (xu2024faithful). More recent work embeds long-form CoT directly into inference, enabling reflection, error correction, and alternative reasoning strategies (jaech2024openai; team2024qwq; guo2025deepseek; team2025kimi; yeo2025demystifying; chen2025towards). In this work, we investigate whether CoT reflects genuine reasoning or merely pattern interpolation.
@@ 2.2 Discussion on Illusion of LLM Reasoning
Recent work questions the robustness and faithfulness of these gains (stechly2024chain). A prominent line of research shows that CoT reasoning is highly fragile: semantically irrelevant perturbations, such as distractor phrases or altered symbolic representations, can substantially degrade performance (mirzadehgsm; tang2023large). Other studies find that models favor surface-level reasoning patterns over logical validity. Moreover, reasoning performance scales poorly with task difficulty, with models over-elaborating on simple problems and failing on harder ones (shojaee2025illusion). Concerns about faithfulness further arise from intervention-based analyses showing that final answers often remain unchanged when intermediate steps are corrupted or removed (lanham2023measuring), an effect referred to as the illusion of transparency (chen2025reasoning; bentham2024chainofthought). The opposing perspectives on CoT reasoning call for a systematic understanding of why and when CoT reasoning succeeds or fails.
@@ 2.3 OOD Generalization of LLMs
Out-of-distribution (OOD) generalization remains a central challenge in machine learning (yang2024generalized; yang2023out; budnikov2025generalization; zhang2024can). Prior work shows that pre-trained models face challenges in adapting to new settings when prompted to learn novel functions (wang2024can; garg2022can). Researchers reveal CoT prompting can partially improve OOD generalization ability, especially for tasks that require long reasoning (yao2025unveiling; shen2025codi). However, other work claims such gains are not intrinsic. For example, strong arithmetic generalization emerges only when algorithmic biases are encoded in positional representations (cho2024position), and finer-grained CoT supervision during training substantially improves OOD performance (wang2025chain). Recent studies further indicate that LLM generalizes reliably when common latent structures are shared across distributions (wang2025theoretical; li2025training). In light of these brilliant findings, we propose rethinking CoT reasoning through a data distribution lens, dissecting CoT reasoning into task, length, and format, and systematically investigating through controlled experiments. We further provide a Comparison with representative work in Appendix A.4.
3 The Proposed Data Distribution Lens
We propose the data distribution lens to understand why and when CoT reasoning succeeds or fails. We hypothesize that CoT reasoning reflects a structured inductive bias learned from in-distribution data, enabling models to conditionally generate reasoning trajectories that approximate those observed during training. As such, the effectiveness of CoT reasoning is fundamentally governed by the nature and degree of distribution discrepancy between training data and test queries (rather than by model architecture or scale).
@@ 3 The Proposed Data Distribution Lens
To formalize this view, we first introduce notation for the training and test distributions. Let $\mathcal{D}_{\text{train}}$ denote the training distribution over input-output pairs $(x,y)$, where $x$ represents a reasoning problem and $y$ denotes the solution sequence (including intermediate reasoning traces). During training, the model learns a parametric mapping $f_{\theta}(x)\approx y$ by minimizing the empirical training risk
$\hat{R}_{\text{train}}(f_{\theta})=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl(f_{\theta}(x_{i}),y_{i}\bigr)$ (1)
where $(x_{i},y_{i})\sim\mathcal{D}_{\text{train}}$ are i.i.d. samples and $\ell$ is a loss function (e.g., cross-entropy). The corresponding expected (population) training risk is
$R_{\text{train}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{train}}}\bigl[\ell(f_{\theta}(x),y)\bigr]$ (2)
@@
At inference time, given a test query sampled from a potentially different distribution $\mathcal{D}_{\text{test}}$, the model generates a response. The expected test risk is
$R_{\text{test}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}\bigl[\ell(f_{\theta}(x),y)\bigr]$ (3)
Definition 3.1 (Distribution Discrepancy).
@@
Given training distribution $\mathcal{D}_{\text{train}}$ and test distribution $\mathcal{D}_{\text{test}}$, we define the distribution discrepancy as
$\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}):=\operatorname{TV}(\mathcal{D}_{\text{train}}\mathcal{D}_{\text{test}})$ (4)
where $\operatorname{TV}(P,Q)$ is the total variation distance,
$\operatorname{TV}(P,Q):=\sup_{A}|P(A)-Q(A)|=\frac{1}{2}\int|dP-dQ|$ (5)
Theorem 3.1 (Generalization Bound).
Assume the loss is bounded, i.e., for all $(x,y)$, $0\leq\ell(f_{\theta}(x),y)\leq B$. Let $\{(x_{i},y_{i})\}_{i=1}^{n}$ be i.i.d. samples from $\mathcal{D}_{\text{train}}$ and let $\hat{R}_{\text{train}}(f_{\theta})$ be the empirical training risk defined above. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the draw of the training sample, the expected test risk satisfies
@@
$R_{\text{test}}(f_{\theta})\!\leq\!\hat{R}_{\text{train}}(f_{\theta})\!+\!2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})\!+\!B\sqrt{\frac{\log(1/\delta)}{2n}}.$ (6)
The proof is provided in Appendix H.1.
Theorem 3.1 provided a theoretical foundation for the data distribution lens. Guided by it, we identify three critical dimensions along which distribution shifts can occur: task, length, and format.
$\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})=\Phi\left(\Delta_{\text{task}},\Delta_{\text{length}},\Delta_{\text{format}}\right)$ (7)
where $\Phi$ is a monotonically increasing composition function that aggregates all discrepancies. $\Delta_{\text{task}}$, $\Delta_{\text{length}}$, and $\Delta_{\text{format}}$ measure the distribution discrepancy introduced by unseen tasks, various lengths, and prompt format variants.
Figure 2: Framework of DataAlchemy. DataAlchemy provides an abstract representation system that distills various real-world NLP tasks into key components: atoms, elements, and transformations. By varying these components, we curate data that exhibits various distribution discrepancies following task, length, and format generalization. DataAlchemy achieves full and fine-grained control over the entire evaluation pipeline. Later, we train models from scratch to avoid data leakage and employ controlled experiments to rigorously test the hypotheses.
4 DataAlchemy: A Controllable Environment
To empirically validate the data distribution lens, we introduce DataAlchemy, an abstract, controllable, and clean environment. It distills real-world NLP tasks into basic atoms, elements, and transformations as illustrated in Figure 2.